# Chain-of-Thought Reasoning Improves Context-Aware Translation with Large Language Models

### v1

@@ A.3.  Results: Lexical Choice
A.3.1.  Accuracy and Inconsistency Scores, Elapsed Time and Cost
This section reports results on the contrastive lexical choice benchmark, for each LLM and prompt configuration.
Model (none, simple) (simple, simple) (detailed, simple) (simple, step-by-step)
@@ A.3.1.  Accuracy and Inconsistency Scores, Elapsed Time and Cost
Table 5: Accuracy (acc $\uparrow$) and inconsistency (inc $\downarrow$) across models and prompts on the contrastive lexical choice validation set (separated by ‘/’ in each cell). The systems are ranked by decreasing accuracy of the best overall prompt, which is the (simple, step-by-step) prompt.
Model (none, simple) (simple, simple) (detailed, simple) (simple, step-by-step)
@@
Model (none, simple) (simple, simple) (detailed, simple) (simple, step-by-step)
@@
Table 7: Total API cost (USD) by prompt configuration for OpenAI models on the lexical choice validation set. Experiments were run in spring 2025; reported costs reflect that period’s pricing and our token usage.
@@
A.3.2.  Graphical Representation of Scores
Figure 6: Mean accuracy (top) and inconsistency (bottom; lower is better) across prompt configurations for each model on the lexical choice validation set.
@@ B.1.  Prompt Variants
This section presents the full text of the prompts used in our generative evaluation. Similarly to the contrastive prompts, each configuration has a system prompt and a user prompt. and the placeholders appear in curly braces: {context_en}, {source_sentence}, {context_fr}, {option1}, {option2}. In the generative setting, the model must produce the French translation of the second sentence, with output format rules depending on the reasoning prompt used.

### v2

@@ A.3.  Results: Lexical Choice
Below are the results on the contrastive lexical choice benchmark, for each LLM and prompt type, in terms of accuracy and inconsistency scores, elapsed time and cost of LLM.
System and user prompts
Model None & simple Both simple Detailed & simple) Simple & step-by-step
@@
Table 11: Accuracy ($\uparrow$) and inconsistency ($\downarrow$) across models and prompts on the contrastive lexical choice validation set (separated by ‘/’ in each cell). The systems are ranked by decreasing accuracy of the best overall prompt, which is the fourth one.
System and user prompts
Model None & simple Both simple Detailed & simple) Simple & step-by-step
@@
System and user prompts
Model None & simple Both simple Detailed & simple) Simple & step-by-step
@@
Table 13: Total API cost (USD) by prompt configuration for OpenAI models on the lexical choice validation set. Experiments were run in summer 2025; reported costs reflect that period’s pricing and our token usage.
@@
A.3.1.  Graphical Representation of Scores
@@ B.1.  Prompt Variants
This section presents the full text of the prompts used in our generative evaluation (translation task). Similarly to the contrastive prompts, each configuration has a system prompt and a user prompt. and the placeholders appear in curly braces: {context_en}, {source_sentence}, {context_fr}, {option1}, {option2}. In the generative setting, the model must produce the French translation of the second sentence, with output format rules depending on the reasoning prompt used.